OpenStudy (anonymous):

Identify the correct x-intercept and y-intercept of the equation 15x - 5y = 30. x-intercept = (-2, 0); y-intercept = (0, 6) x-intercept = (2, 0); y-intercept = (0, -6) x-intercept = (0, -6); y-intercept = (2, 0) x-intercept = (0, 6); y-intercept = (-2, 0)

OpenStudy (ivettef365):

convert as y = mx + b, the b is the y-int so you have 15x -5y = 30 -5y = -15x + 30 --- --- --- -5 -5 -5

OpenStudy (ivettef365):

then you have y = 3x - 6, so (0,-6) is the y-int to find the x, then find out what is the value of y when x =0

OpenStudy (ivettef365):

I meant when y = 0, then you have 15x - 15(0) = 30 15x = 30 x = 2, so it is (2,0)

OpenStudy (anonymous):

thank you so much and thanks for not just giving a answer and actually explaining it.

OpenStudy (ivettef365):

yw :)